# Asymptotically Unbiased Synthetic Control Methods by Distribution Matching

### v3

@@ Abstract
Synthetic Control Methods (SCMs) have become an essential tool for comparative case studies. The fundamental idea of SCMs is to estimate the counterfactual outcomes of a treated unit using a weighted sum of the observed outcomes of untreated units. The accuracy of the synthetic control (SC) is critical for evaluating the treatment effect of a policy intervention; therefore, the estimation of SC weights has been the focus of extensive research. In this study, we first point out that existing SCMs suffer from an endogeneity problem, the correlation between the outcomes of untreated units and the error term of the synthetic control, which yields a bias in the treatment effect estimator. We then propose a novel SCM based on density matching, assuming that the density of outcomes of the treated unit can be approximated by a weighted average of the joint density of untreated units (i.e., a mixture model). Based on this assumption, we estimate SC weights by matching the moments of treated outcomes with the weighted sum of moments of untreated outcomes. Our proposed method has three advantages over existing methods: first, our estimator is asymptotically unbiased under the assumption of the mixture model; second, due to the asymptotic unbiasedness, we can reduce the mean squared error in counterfactual predictions; third, our method generates full densities of the treatment effect, not merely expected values, which broadens the applicability of SCMs. We provide experimental results to demonstrate the effectiveness of our proposed method. [^1] This study was presented at the Workshop on Counterfactuals in Minds and Machines at the International Conference on Machine Learning in July 2023 and at the International Conference on Econometrics and Statistics in August 2023.
@@ 1 Introduction
Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have gained attention as a crucial tool for program evaluation in comparative case studies across fields such as economics, statistics, and machine learning. According to Athey and Imbens (2017), SCMs are “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
@@
SCMs posit a scenario in which multiple units exist, with one unit subjected to a policy intervention (treated unit) at a certain time period, while the rest remain untreated. The focus is on estimating the treatment effect of the treated unit, i.e., the difference between its observed factual outcome and unobserved counterfactual outcome. Untreated units are used to account for unobserved trends in the outcome over time that are unrelated to the effect of the policy intervention. The key insight is that an optimally weighted average of the untreated units, referred to as the synthetic control (SC) unit, provides a more appropriate estimate of the counterfactual than estimating the counterfactual using only the treated unit. See Figure 1 for an illustration. SCMs have been applied in a variety of settings, such as studying the effect of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017).
@@
For instance, Abadie et al. (2010) used SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988. Smoking rates in California decreased following the program’s implementation, but it was unclear if this decrease was a result of the program or a long-term trend. To investigate the treatment effect, annual per-capita cigarette sales data across several states that did not implement a similar tobacco control program were used as untreated units. They then applied SCM to estimate California’s counterfactual outcome: the smoking rate in California in 1989 had the tobacco control program not been implemented. For further details, also see Abadie (2021).
A significant challenge in SCMs is the typical assumption that the expected outcome of the treated unit can be represented by the weighted sum of untreated units. However, Ferman and Pinto (2021) points out that there exists a non-negligible asymptotic bias in the obtained estimator. According to Ferman and Pinto (2021), this asymptotic bias is caused by a kind of endogeneity in the linear regression model deduced from the assumption, due to a correlation between the outcomes of untreated units and the error term of the linear model. This problem is akin to endogeneity caused by measurement error (Greene, 2003). We refer to this endogeneity as implicit endogeneity.
To address the aforementioned issue, we propose to assume a mixture model instead of a linear relationship among expected outcomes. Specifically, we assume that the density of the outcome of the treated unit can be approximated by a linear combination of those of the untreated units. This mixture model assumption also implies that each moment of the outcome of the treated unit can be represented by a linear combination of those of the untreated units. Therefore, we can estimate the SC weights by matching all moments between the outcome of the treated unit and those of the untreated units. We refer to our method as the Density Matching SCM (DMSCM). Unlike existing methods, our proposed estimator is asymptotically unbiased. We validate the efficacy of our proposed method through simulation studies and empirical analysis.
Our proposed method corresponds to a simplification of the distributional SCMs by Gunsilius (2023). Although Gunsilius (2023) does not explicitly point out the implicit endogeneity problem, it observes that their result does not coincide with the result under the standard SCMs. This is because under the mixture model assumption, while Gunsilius (2023) returns a consistent estimator for the treatment effect, existing methods do not. We connect the distributional SCMs to the asymptotic bias problem and propose a simplified estimation method for SCMs by distribution matching.
In summary, our contributions are as follows:
  • Reconsidering the assumptions necessary for treatment effect estimation in SCMs.
  • Proposing a density-matching method for asymptotically unbiased SCMs.
  • Simplifying the density-matching method.
1.1 Related Work
@@
Following the pioneering work by Abadie and Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015). These SCMs, including their variants such as those by Doudchenko and Imbens (2016) and Ben-Michael et al. (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights.
Gunsilius (2023) explores distribution SCMs (DiSCo), which estimate SC weights to match the distribution of the outcomes of the treated unit and the weighted sum of the distributions of the untreated units. [^2] The method of Gunsilius (2023) is called DiSCo by Dijcke (2024). His method is based on minimizing the Wasserstein distance, which can be simplified as matching the quantiles in the setting of SCMs. Their assumption about identification can be interpreted as a variant of mixture models.
@@
From a theoretical viewpoint, Abadie et al. (2010) demonstrate that, if there are weights that provide a perfect pre-treatment fit, then the bias of the SC estimator is bounded by a function that converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases, even when the number of untreated units $J$ is fixed. However, Ferman and Pinto (2021) finds that if the pre-treatment fit is imperfect, then the conventional SCMs do not yield a consistent estimator of the treatment effect due to the bias caused by the correlation between the outcomes of the untreated units and the error term in the posited linear models. We review this problem in Section 4 and name it as the implicit endogeneity.
To address this bias issue, existing studies have proposed several approaches. Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where the number of untreated units increases ($J\to\infty$) as well as the length of the pre-intervention periods $T_{0}$ also increases.
The use of mixture models has a close relationship with linear factor models, assumed in existing studies (Abadie, 2021). We justify the use of the former by employing arguments from Shi et al. (2022) and Nazaret et al. (2023). Our proposed method can be considered a new interpretation of existing SCMs. By reinterpreting existing SCMs through density matching, in addition to addressing the implicit endogeneity problem, we address two problems in the SCM literature. First, our proposed method improves the pre-treatment fit by debiasing the estimator. Second, our method generates full counterfactual densities of the treated unit, rather than merely the expected value. As a result, we can conduct statistical inference on the distribution, or any function of the distribution, such as quantiles (Gunsilius, 2023), Lorenz curves (Gastwirth, 1971), and expected social welfare (Abadie, 2002).
An excellent illustration of the relevance of estimating distribution functions is Ropponen (2011), who applies the changes-in-changes estimator by Athey and Imbens (2006) to estimate the effect of minimum wage changes on employment levels in order to reconcile the classical conflicting results about the minimum wage debate in Card and Krueger (2000) and Neumark and Wascher (2000). Spiess et al. (2023) investigates high-dimensional linear regression models in SCMs.
The exploration of identification assumptions has been attempted in the existing literature. Ferman and Pinto (2021) conducts a comprehensive investigation to discern the conditions under which SCMs can accurately predict counterfactual outcomes.
This study also answers an important open question raised by Gunsilius (2023). Gunsilius (2023) discusses that an optimal weight obtained by the density matching differs from that obtained by the classical least squares approach. Gunsilius (2023) conjectures that the classical least squares approach estimates a weight to match the first moment, while the density matching approach estimates a weight to match all moments. However, we point out that the classical least squares approach estimates a weight not only to match the first moment but also tends to incur bias due to the implicit endogeneity.
Concerning the assumptions about SCMs, Wan et al. (2018) points out that the SCMs essentially assume mixture models, which distinguish SCMs from traditional panel data analysis.
On another note, Shi et al. (2022) introduces fine-grained models to investigate hidden assumptions inherent to SCMs, which also explores a connection between the linear combination of expected outcomes and the linear factor models. Nazaret et al. (2023) extends these findings into the realm of model misspecification. Their formulation is also based on the mixture models.
@@
Independently of us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias pointed out by Ferman and Pinto (2021). [^3] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and we and he conducted our studies completely independently. While our and his formulations about the endogeneity are similar, the approaches are significantly different. We address the problem from the viewpoint of density matching, which is an approach similar to Gunsilius (2023). In contrast, Fry (2024) uses an instrumental variable method by employing outcomes of untreated units that are not used in SCMs as instrumental variables.
@@
1.2 Organizations
In Section 2, we formulate the problem. In Section 3, we review an idea of SCMs. In Section 4, we point out the problem of implicit endogeneity. Then, in Section 5, we propose our method under the assumption of mixture models. In Section 6, we investiate the asymptotic property of our proposed method. In Section 7, we present a statistical inference method for our proposed method. In Section 8, we discuss related topics and existing studies. We numerically investigate the performance of the proposed estimators through extensive simulations in Section 9.
@@ 2 Problem Setting
Suppose that there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$ where $3\leq T\in\mathbb{N}$. Without loss of generality, we assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest. A set of potential comparisons, $j\in\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consists of untreated units not affected by the intervention. We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, such that $2\leq T_{0}<T$. Observations at period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations at period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention. Let $T_{1}\coloneqq|\mathcal{T}_{1}|=T-T_{0}$.
@@ 2.1 Potential Outcomes
First, following the Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974), we define the potential outcomes of the policy intervention. For each unit $j\in\mathcal{J}$ and period $t\in\mathcal{T}$, there are two outcomes of interest, $\left(Y^{I}_{j,t},Y^{N}_{j,t}\right)\in\mathbb{R}^{2}$, corresponding to scenarios with and without the intervention. For each $j\in\mathcal{J}$ and $t\in\mathcal{T}$, let $F^{I}_{j,t}(y)$ and $F^{N}_{j,t}(y)$ be the cumulative distribution functions (CDFs) for $Y^{I}_{j,t}$ and $Y^{N}_{j,t}$, respectively.
@@ 2.2 Observations
We then define our observations. For each unit $j\in\mathcal{J}$, we can only observe one of the outcomes corresponding to the actual intervention; that is, we observe $Y_{j,t}\in\mathbb{R}$ such that
$Y_{0,t}=\begin{cases}Y^{N}_{0,t}&\forall t\in\mathcal{T}_{0}\\ Y^{I}_{0,t}&\forall t\in\mathcal{T}_{1}\end{cases},\qquad\mathrm{and}\qquad Y_{j,t}=Y^{N}_{j,t}\qquad\forall j\in\mathcal{J}^{U},\forall t\in\mathcal{T}.$
We aim to estimate the effect of the policy intervention by predicting $Y^{I}_{0,t}$ from $(Y^{N}_{j,t})_{j\in\mathcal{J}^{U}}$ and by comparing this prediction to the observed $Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$. We will define the treatment effect more formally in the following section.
@@ 2.3 Treatment Effect
In SCMs, there are several ways to define the treatment effect $\tau_{0,t}\in\mathbb{R}$ for each $t\in\mathcal{T}_{1}$ of the intervention for the treated unit. First, for each $t\in\mathcal{T}_{1}$, we define the treatment effect on the treated unit $j=0$ as
$\tau_{0,t}\coloneqq\mathbb{E}\big{[}Y^{I}_{0,t}\big{]}-\mathbb{E}\big{[}Y^{N}_{0,t}\big{]}.$
@@
$Y^{I}_{j,t}=\tau_{j,t}+Y^{N}_{j,t}.$
where $\delta_{t}$ is an unobserved common factor with constant factor loadings across units, $c_{j}$ is an unknown time-invariant fixed effect, $\lambda_{t}$ is a $(1\times F)$ vector of unobserved common factors, $\mu_{j}$ is a $(F\times 1)$ vector of unknown factor loadings, and the error terms $\epsilon_{j,t}$ are unobserved idiosyncratic shocks. Here, while $c_{j}$, $\mu_{j}$, and $\tau_{0,t}$ are non-random variables, $\delta_{t}$, $\lambda_{t}$, and $\epsilon_{j,t}$ are random variables.
Under the linear factor model, the treatment effect is given by $\tau_{0,t}=Y^{I}_{0,t}-Y^{N}_{0,t}$, a non-random variable by definition of the factor model. Therefore, the expression $\tau_{0,t}\coloneqq\mathbb{E}\big{[}Y^{I}_{0,t}\big{]}-\mathbb{E}\big{[}Y^{N}_{0,t}\big{]}$ also holds under the factor model.
Comparative case studies aim to evaluate $\tau_{0,t}$ by predicting $Y^{N}_{0,t}$ for $t>T_{0}$, which is the counterfactual outcome that would have been observed for the treated unit in the absence of the intervention. By definition, $Y^{N}_{0,t}$ is a counterfactual outcome because of the intervention after $t=T_{0}$. We usually use one unit or a small number of units that have similar characteristics to the treated unit at the time of the intervention but are not exposed to the intervention. However, when the data consist of a few units, such as regions or countries, it is often difficult to find a single untreated unit that provides a suitable comparison for the unit treated by the policy intervention of interest.
Distributional Treatment Effects. In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as the quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002). In this study, we consider estimating $p^{I}_{0,t}(y)$ and $p^{N}_{0,t}(y)$ for $t\in\mathcal{T}_{1}$, where outcomes corresponding to $p^{I}_{0,t}(y)$ are unobservable. The treatment effect is then captured by $p^{I}_{0,t}(y)$ and $p^{N}_{0,t}(y)$, or functionals of these distributions, such as QTEs and expected social welfare. Depending on the application, we may adjust the target density, for instance, $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$. We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus and Oprescu, 2022; Chikahara et al., 2022). In the context of SCMs, Gunsilius (2023) and Chen (2020) have estimated the treatment effect.

### v4

# Asymptotically Unbiased Synthetic Control Methods by Density Matching

@@ Abstract
Synthetic Control Methods (SCMs) have become a fundamental tool for comparative case studies. The core idea behind SCMs is to estimate treatment effects by predicting counterfactual outcomes for a treated unit using a weighted combination of observed outcomes from untreated units. The accuracy of these predictions is crucial for evaluating the treatment effect of a policy intervention. Subsequent research has therefore focused on estimating SC weights. In this study, we highlight a key endogeneity issue in existing SCMs—namely, the correlation between the outcomes of untreated units and the error term of the synthetic control, which leads to bias in both counterfactual outcome prediction and treatment effect estimation. To address this issue, we propose a novel SCM based on density matching, assuming that the outcome density of the treated unit can be approximated by a weighted mixture of the joint density of untreated units. Under this assumption, we estimate SC weights by matching the moments of the treated outcomes with the weighted sum of the moments of the untreated outcomes. Our method offers three advantages: first, under the mixture model assumption, our estimator is asymptotically unbiased; second, this asymptotic unbiasedness reduces the mean squared error in counterfactual predictions; and third, our method provides full densities of the treatment effect rather than just expected values, thereby broadening the applicability of SCMs. Finally, we present experimental results that demonstrate the effectiveness of our approach. [^1] This study was presented at the Workshop on Counterfactuals in Minds and Machines at the International Conference on Machine Learning in July 2023 and at the International Conference on Econometrics and Statistics in August 2023.
@@ 1 Introduction
Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have emerged as a crucial tool for program evaluation in comparative case studies across disciplines such as economics, statistics, and machine learning. As noted by Athey and Imbens (2017), SCMs represent “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
@@
In SCMs, there are multiple units, where one unit receives a policy intervention (treated unit) at a certain time while the remaining units remain untreated. The objective is to estimate the treatment effect for the treated unit, defined as the difference between its observed factual outcome and its unobserved counterfactual outcome. Untreated units help account for unobserved trends in the outcome over time that are unrelated to the policy intervention. The key idea is that an optimally weighted combination of the untreated units, referred to as the synthetic control (SC) unit, serves as an appropriate estimate of the counterfactual outcome. See Figure 1 for an illustration. d SCMs have been widely applied in various empirical studies, such as evaluating the impact of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017). For instance, Abadie et al. (2010) applies SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988. While smoking rates in California declined following the program’s implementation, it remained unclear whether this decrease was attributable to the policy. For estimating the treatment effect, annual per-capita cigarette sales data from states that did not implement a similar tobacco control program were used as untreated units. SCM was then applied to construct California’s counterfactual outcome, representing the smoking rate in the absence of the tobacco control program.
A key challenge in SCMs is the assumption, made in most existing studies (Abadie, 2002), that the expected outcome of the treated unit can be represented as a weighted sum of the outcomes of untreated units. Ferman and Pinto (2021) highlights that this assumption leads to a non-negligible bias in the estimator. According to Ferman and Pinto (2021), this bias arises from endogeneity in the linear regression model implied by the SCM assumption due to the correlation between the outcomes of untreated units and the error term. This issue is analogous to the endogeneity caused by measurement error (Greene, 2003). We refer to this problem as implicit endogeneity.
To address this issue, we propose an alternative approach based on a mixture model assumption rather than a linear relationship among expected outcomes. Specifically, we assume that the density of the treated unit’s outcome can be approximated by a weighted combination of the densities of the untreated units. This assumption further implies that each moment of the treated unit’s outcome can be represented as a weighted sum of the moments of the untreated units. Leveraging this insight, we estimate SC weights by matching the moments of the treated unit’s outcome to those of the untreated units. We refer to our method as the Density Matching SCM (DMSCM). Unlike existing methods, our proposed estimator is asymptotically unbiased if the mixture model assumption is correct. We validate the effectiveness of our approach through both simulation studies and empirical analysis.
The mixture model assumption has appeared in previous and concurrent studies, including Wan et al. (2018), Gunsilius (2023), and Shi et al. (2022). Our proposed method can be viewed as a simplification of the distributional SCMs introduced by Gunsilius (2023). While Gunsilius (2023) does not explicitly address the implicit endogeneity problem, they note that their results diverge from those obtained under standard SCMs. This discrepancy arises because, under the mixture model assumption, Gunsilius (2023) provides a consistent estimator for the treatment effect, whereas existing SCM methods do not. We establish a connection between distributional SCMs and the asymptotic bias problem and propose a simplified estimation method based on density matching.
In summary, our contributions are as follows:
  • Reexamining the assumptions necessary for treatment effect estimation in SCMs.
  • Proposing a density-matching approach for asymptotically unbiased SCMs.
  • Simplifying the implementation of density matching in SCMs.
We formulate the problem in Section 2. Next, in Section 3, we review the basic idea of SCMs. In Section 4, we highlight the problem of implicit endogeneity. In Section 5, we then propose our method under the mixture model assumption. We investigate the asymptotic properties of this method in Section 6 and present a statistical inference method in Section 7. In Section 8, we discuss related topics and existing studies. Finally, in Section 9, we examine the performance of our estimators through extensive simulations.
@@
Related work. Following the pioneering work by Abadie and Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015). These SCMs, including their variants such as those by Doudchenko and Imbens (2016) and Ben-Michael et al. (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights. We refer to such SCMs as Least-Squares SCMs (LS-SCMs).
The bias of the LS-SCMs strongly depends on whether there exist weights that can predict the outcome of the treated unit without error. In other words, a random outcome of the treated unit must be replicable by the sum of outcomes of untreated units. Since such weights yield a perfect fit in the estimation process using observations from the pre-treatment period, this situation is called pretreatment-fit.
@@
Abadie et al. (2010) demonstrates that under a pre-treatment fit, the bias of the LS-SCM estimator converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases. However, Ferman and Pinto (2021) finds that if the pre-treatment fit does not hold, then the LS-SCM estimator has a bias and is not consistent. The cause of this bias is the correlation between the outcomes of the untreated units and the error term in the posited linear model, which we refer to as implicit endogeneity.
@@
Several methods have been proposed to address this bias issue. Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where both the number of untreated units ($J\to\infty$) and the length of the pre-intervention periods ($T_{0}$) grow large. Concurrently with us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias reported by Ferman and Pinto (2021). [^2] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and he and we conducted our studies completely independently. While our formulation and his regarding endogeneity are similar, our approach differs significantly. We address the problem through density matching, similar to the strategy in Gunsilius (2023), whereas Fry (2024) employs an instrumental variable approach using outcomes of untreated units not used in SCMs as instruments.
In this study, we focus on mixture models among outcome distributions, an assumption also employed in Wan et al. (2018), Gunsilius (2023), Shi et al. (2022), and Nazaret et al. (2023). Although these studies introduce mixture models with motivations other than the bias problem, we highlight that such models can effectively circumvent bias under imperfect pre-treatment fit.
Here, we briefly review existing work using mixture models. Wan et al. (2018) points out that SCMs essentially rely on mixture models, distinguishing them from traditional panel data analysis. Gunsilius (2023) proposes distribution SCMs (DiSCo), which estimate SC weights by matching the outcome distributions of the treated unit and the weighted sum of the distributions of the untreated units. Shi et al. (2022) introduces fine-grained models to elucidate hidden assumptions in SCMs and explores connections to linear factor models. Nazaret et al. (2023) extends these insights by analyzing model misspecification under a mixture model formulation.
This study also addresses a question raised by Gunsilius (2023). While Gunsilius (2023) suggests that standard least squares SCMs try to match the first moment of the outcome distribution (and therefore differ from distributional approaches), we stress that least squares approaches are additionally vulnerable to implicit endogeneity, which can generate bias. Our density-matching approach avoids this problem by matching a broader set of distributional features.
@@ 2 Problem Setting
Suppose there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$, where $3\leq T\in\mathbb{N}$. Without loss of generality, assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest. The set of potential control units is $\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consisting of units not affected by the intervention. We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, with $2\leq T_{0}<T$. Observations in period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations in period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention. Let $T_{1}\coloneqq|\mathcal{T}_{1}|=T-T_{0}$.
@@ 2.1 Potential Outcomes
Following the Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974), we define the potential outcomes of the policy intervention. For each unit $j\in\mathcal{J}$ and period $t\in\mathcal{T}$, there are two outcomes of interest, $\left(Y^{I}_{j,t},Y^{N}_{j,t}\right)\in\mathbb{R}^{2}$, corresponding to scenarios with and without the intervention. For each $j\in\mathcal{J}$ and $t\in\mathcal{T}$, let $F^{I}_{j,t}(y)$ and $F^{N}_{j,t}(y)$ be the cumulative distribution functions (CDFs) for $Y^{I}_{j,t}$ and $Y^{N}_{j,t}$, respectively.
@@ 2.2 Observations
We next define our observations. For each unit $j\in\mathcal{J}$, only one of the two potential outcomes can be observed, corresponding to what actually happened. Specifically, for each $j\in\mathcal{J}$, we observe $Y_{j,t}\in\mathbb{R}$ such that
$Y_{0,t}=\begin{cases}Y^{N}_{0,t}&\forall t\in\mathcal{T}_{0},\\ Y^{I}_{0,t}&\forall t\in\mathcal{T}_{1},\end{cases}\quad\mathrm{and}\quad Y_{j,t}=Y^{N}_{j,t}\quad\forall j\in\mathcal{J}^{U},\forall t\in\mathcal{T}.$
We aim to estimate the effect of the policy intervention by predicting $Y^{I}_{0,t}$ from $\left(Y^{N}_{j,t}\right)_{j\in\mathcal{J}^{U}}$ and comparing that prediction to the observed $Y^{N}_{0,t}$ for $t\in\mathcal{T}_{1}$. We define the treatment effect more formally in the next subsection.
@@ 2.3 Treatment Effect
In SCMs, there are several ways to define the treatment effect $\tau_{0,t}\in\mathbb{R}$ for each $t\in\mathcal{T}_{1}$ (i.e., after the intervention). First, for each $t\in\mathcal{T}_{1}$, we can define the treatment effect for the treated unit $j=0$ as
$\tau_{0,t}\coloneqq\mathbb{E}[Y^{I}_{0,t}]-\mathbb{E}[Y^{N}_{0,t}].$
@@
$Y^{I}_{j,t}=\tau_{j,t}+Y^{N}_{j,t},$
where $\delta_{t}$ is an unobserved common factor with constant factor loadings across units, $c_{j}$ is an unknown time-invariant fixed effect, $\lambda_{t}$ is a $(1\times F)$ vector of unobserved common factors, $\mu_{j}$ is a $(F\times 1)$ vector of unknown factor loadings, and $\epsilon_{j,t}$ is an unobserved idiosyncratic shock. Here, $c_{j}$, $\mu_{j}$, and $\tau_{0,t}$ are non-random variables, while $\delta_{t}$, $\lambda_{t}$, and $\epsilon_{j,t}$ are random variables.
Under the linear factor model, the treatment effect is given by $\tau_{0,t}=Y^{I}_{0,t}-Y^{N}_{0,t}$, a non-random variable by definition. Consequently, the expression $\tau_{0,t}\coloneqq\mathbb{E}[Y^{I}_{0,t}]-\mathbb{E}[Y^{N}_{0,t}]$ also holds under the factor model.
Comparative case studies aim to evaluate $\tau_{0,t}$ by predicting $Y^{N}_{0,t}$ for $t>T_{0}$, the counterfactual outcome that would have been observed for the treated unit had the intervention not occurred. Because the intervention takes place after $t=T_{0}$, $Y^{N}_{0,t}$ is a counterfactual outcome. In settings where data consist of only a few units (e.g., regions or countries), it can be difficult to find a single untreated unit that provides a suitable comparison. For this problem, SCMs enable us to construct a counterfactual outcome using a predictor based on the outcomes of the untreated units. Under suitable model assumptions that relate $Y^{N}_{0,t}$ to $\left(Y^{N}_{j,t}\right)_{j\in\mathcal{J}^{U}}$, we can construct the predictor using observations from the pre-intervention period and then predict the unobserved counterfactual outcome in the post-intervention period. Consequently, we estimate the treatment effect by comparing the predicted $Y^{N}_{0,t}$ with the observed $Y^{I}_{0,t}$.
Distributional Treatment Effects. In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002). By applying our method, we can also estimate the distribution of the counterfactual outcome, $p^{N}_{0,t}(y)$, for $t\in\mathcal{T}_{1}$. The treatment effect is then captured by the difference between $p^{I}_{0,t}(y)$ (the distribution of the actual outcome) and $p^{N}_{0,t}(y)$, or by functionals of these distributions, such as QTEs and expected social welfare. Depending on the application, we may also adjust the target density, for example, by averaging across all post-intervention periods: $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$. We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus and Oprescu, 2022; Chikahara et al., 2022). In the context of SCMs, Gunsilius (2023) and Chen (2020) have proposed methods for estimating such distributional treatment effects.